OpenStudy (anonymous):

find the domain and range of each function. (a)F(x)=(x^2-4)^1/2 (b)G(x)=(x^2+9)^1/2 (c)H(x)=1/x^2 (D)t(x)=cos4x (e)H(G(x))

OpenStudy (anonymous):

\[ f(x)=\sqrt{x^2-4}\] you have to make sure that \[x^2-4\geq 0\] and since \[x^2-4\] is a parabola facing up it is positive outside the zeros and negative between them so your domain is \[(-\infty,-2]\cup [2,\infty)\]

OpenStudy (anonymous):

range is \[y\geq 0\] because the radical sign means positive root

OpenStudy (anonymous):

\[ g(x)=\sqrt{x^2+9}\] and now \[x^2+9\geq 9\] for all x, so it is never negative and therefore domain is all real numbers. range is \[y\geq 0\] for the same reason as the other one

OpenStudy (anonymous):

\[H(x)=\frac{1}{x^2}\] domain is all numbers except 0 because you cannot divide by zero. range is \[y>0\] since a perfect square is always greater than or equal to zero and this is never zero because the numerator is 1

OpenStudy (anonymous):

cosine has domain all real numbers. range is \[[-1,1]\]

OpenStudy (anonymous):

THANK YOU!

OpenStudy (anonymous):

yw

OpenStudy (anonymous):

but how would i do (d) and (e) ?